OpenStudy (anonymous):

what is c when 2/5=4/5c? And how is that solved?

OpenStudy (anonymous):

Cross multiply... 2*5c=5*4 10c=20 c=?

OpenStudy (anonymous):

CalebBeavers, thanks for the reply, but these are fractions I'm working with. 2/5=4/5c

OpenStudy (anonymous):

Yeah, you cross multiply....

OpenStudy (anonymous):

5*4=2*5c

OpenStudy (anonymous):

Now I get it. Thanks for the help.

OpenStudy (anonymous):

If c is 2, then 4/5 of 2 = 1.6, and 2/5 = .4, right?